# StoryTeller: Improving Long Video Description through Global Audio-Visual Character Identification

### v1

@@ 6 Conclusion
In this paper, we figure out audio-visual character identification as a crucial step for generating detailed and consistent descriptions of long videos. Based on this insight, we propose StoryTeller, a system designed to generate dense, plot-level descriptions for long videos. The system is composed of three main modules: video segmentation, audio-visual character identification, and description generation. We also introduce a new dataset, MovieStory101, which includes 5,982 three-minute movie clips from 187 movies. To effectively and efficiently evaluate long video descriptions, we present MovieQA, a large-scale multiple-choice question-answering dataset aligned with the MovieStory101 test set. We assess descriptions by inputting them into GPT-4 to answer these questions, using accuracy as an automatic evaluation metric. Experiments demonstrate that StoryTeller outperforms all open-source and closed-source baselines on MovieQA. Specifically, StoryTeller surpasses Gemini-1.5-pro by 9.5% in accuracy on MovieQA. In human side-by-side evaluations, StoryTeller shows a +15.56% advantage over Gemini-1.5-pro. Additionally, we verify that incorporating audio-visual character identification significantly enhances video descriptions for all LVLMs. Notably, Gemini-1.5-pro and GPT-4o achieve relative improvements of 5.5% and 13.0% in accuracy on MovieQA, respectively. These results indicate that StoryTeller ’s methodology provides a general approach for improving long video descriptions.
@@ Appendix A Annotation Guidelines for Movie Description
Appendix B Creation of MovieQA
@@ Appendix D Global ID Generation
Global ID Generation. To determine the optimal threshold for clustering, we utilize data from the MovieStory101 development set to create a validation dataset for threshold selection. We pair all utterances, labeling pairs as positive if both utterances were spoken by the same character, and negative otherwise. We evaluate thresholds ranging from 0.5 to 1.0 on this dataset, with the results presented in Table 7. A threshold of 0.85, which achieved the highest accuracy, was selected. In this context, high precision in identifying whether two utterances were spoken by the same person is crucial, while maintaining a balance with recall. The chosen threshold resulted in a precision of 0.900 and a recall of 0.408, satisfying these requirements.
@@
Table 8: Clustering result with Precision=0.9.

### v2

@@ 6 Conclusion
In this paper, we figure out audio-visual character identification as a crucial step for generating detailed and consistent descriptions of long videos. Based on this insight, we propose StoryTeller, a system for producing dense, plot-level descriptions, composed of three modules: video segmentation, audio-visual character identification, and description generation. We also introduce a new dataset, MovieStory101, with 5,982 three-minute movie clips from 187 movies. To effectively and efficiently evaluate long video descriptions, we present StoryQA, a large-scale multiple-choice question-answering dataset aligned with the MovieStory101 test set. We assess descriptions by inputting them into GPT-4 to answer these questions, using accuracy as an automatic evaluation metric. Experiments demonstrate that StoryTeller outperforms all baselines, surpassing Gemini-1.5-pro by 9.5% in accuracy on StoryQA and achieving a +15.56% advantage in human side-by-side evaluations. Additionally, we verify that incorporating audio-visual character identification significantly enhances video descriptions for all LVLMs, with Gemini-1.5-pro and GPT-4o achieving accuracy improvements of 5.5% and 13.0%, respectively.
@@ Appendix A Annotation Guidelines for Movie Description
Appendix B Creation of StoryQA
@@ Appendix D Global ID Generation
Global ID Generation. To determine the optimal threshold for clustering, we utilize data from the MovieStory101 development set to create a validation dataset for threshold selection. We pair all utterances, labeling pairs as positive if both utterances were spoken by the same character, and negative otherwise. We evaluate thresholds ranging from 0.5 to 1.0 on this dataset, with the results presented in Table 9. A threshold of 0.85, which achieved the highest audio-visual character identification accuracy, was selected. In this context, high precision in identifying whether two utterances were spoken by the same person is crucial, while maintaining a balance with recall. The chosen threshold resulted in a precision of 0.900 and a recall of 0.408, satisfying these requirements.
@@
Table 10: Clustering result with Precision=0.9 on test dataset.